Elimination

5x+y=9
10x-7y=-18

Answer by swincher4391(1107)

Notice 10 = 5*2.
5x + y = 9
10x - 7y = -18

If we multiply the top by -2, then we can add and eliminate x.
-2(5x+y) = -2*9
-10x - 2y = -18
10x - 7y = -18
-9y = -36
y = 4
5x + 4 = 9
5x = 5
x = 1
So our solution is (1,4)